Pleeease please help!!!!
P7E.4 The force constant for the bond in CO is 1857Nm-1. Calculate the vibrational frequencies (in Hz) ofCО 13CО CO, and 13C180. Use integer relative atomic masses for this estimate.
Answer #1

Considering these as simple harmonic oscillators the equation to obtain the vibrational frequency can be written as:

\nu = (1/2\pi )\sqrt{(K/\mu )} .................... Eqn 1

Where \nu is the vibrational frequency in Sec-1 or Hz, K is the force constant in dyne.Cm-1, \mu is the reduced mass in grams.

Reduces mass is given as:

\mu = (m_{1}m_{2})/(m_{1}+m_{2})....................... Eqn 2

Or, \mu = (M_{1}M_{2})/(M_{1}+M_{2})

Here M1 and M2 are the atomic mass in amu and m1 and m2 is the mass of the atoms involved in the bond in grams.

Its given that K = 1857 N.m-1 = (1857\times 105) / 100 dyne.Cm-1 as 1 N = 105 dyne and 1 m = 100 Cm.

So, K = 1857000 dyne.Cm-1

Now Lets calculate the reduced mass using Eqn 2:

For 12C16O:

\mu = (12\times16)/ (12 + 16) = 6.857 amu = 6.857 / (6.023 \times 1023) g = 1.138 \times 10-23g

For 13C16O:

\mu = (13\times16)/ (13 + 16) = 7.1724 amu = 7.1724 / (6.023 \times 1023) g = 1.1908 \times 10-23g

For 12C18O:

\mu = (12\times18)/ (12 + 18) = 7.2 amu = 7.2 / (6.023 \times 1023) g = 1.1954 \times 10-23g

For 13C18O:

\mu = (13\times18)/ (13 + 18) = 7.548 amu = 7.548 / (6.023 \times 1023) g = 1.2532 \times 10-23g

Now substituting these values in Eqn 1 we have:

For 12C16O:

\nu = (1/2\pi )\sqrt{(1857000 dyne.Cm^{-1}/1.138\times 10^{-23}g)}

Or, \nu = (1/2\pi )\sqrt{(1857000 g.Cm.S^{-2}.Cm^{-1}/1.138\times 10^{-23}g)}=6.4324\times 10^{13}S^{-1}orHz

By substituting the values similarly we will get For 13C16O  \nu = 6.2882 \times 1013 Hz,

For 12C18O \nu = 6.2760 \times 1013 Hz and For 13C18O  \nu = 6.1296 \times 1013 Hz